Packaging Solutions Corporation manufactures and sells a wide variety of packaging products. Performance reports are prepared monthly for each department. The planning budget and flexible budget for the Production Department are based on the following formulas, where q is the number of labor-hours worked in a month:

Cost Formulas
Direct labor $16.10q
Indirect labor $4,800 + $1.50q
Utilities $5,500 + $0.70q
Supplies $1,700 + $0.20q
Equipment depreciation $18,600 + $2.70q
Factory rent $8,100
Property taxes $2,600
Factory administration $13,600 + $0.70q

The Production Department planned to work 4,200 labor-hours in March; however, it actually worked 4,000 labor-hours during the month. Its actual costs incurred in March are listed below:

Actual Cost Incurred in March
Direct labor $ 65,940
Indirect labor $ 10,280
Utilities $ 8,830
Supplies $ 2,750
Equipment depreciation $ 29,400
Factory rent $ 8,500
Property taxes $ 2,600
Factory administration $ 15,770

Required:

1. Prepare the Production Department’s planning budget for the month.

2. Prepare the Production Department’s flexible budget for the month.

3. Prepare the Production Department’s flexible budget performance report for March, including both the spending and activity variances.
Answer #1
1
Planning Budget
Budgeted labor hours 4200
Direct labor 67620 =4200*16.1
Indirect labor 11100 =4800+(4200*1.5)
Utilities 8440 =5500+(4200*0.7)
Supplies 2540 =1700+(4200*0.2)
Equipment Depreciation 29940 =18600+(4200*2.7)
Factory rent 8100
Property taxes 2600
Factory administration 16540 =13600+(4200*0.7)
Total expenses 146880
2
Flexible Budget
Actual labor hours 4000
Direct labor 64400 =4000*16.1
Indirect labor 10800 =4800+(4000*1.5)
Utilities 8300 =5500+(4000*0.7)
Supplies 2500 =1700+(4000*0.2)
Equipment Depreciation 29400 =18600+(4000*2.7)
Factory rent 8100
Property taxes 2600
Factory administration 16400 =13600+(4000*0.7)
Total expenses 142500
3
Actual results Spending variance Flexible budget Activity variance Planning budget
Labor hours 4000 4000 4200
Direct labor 65940 1540 U 64400 3220 F 67620
Indirect labor 10280 520 F 10800 300 F 11100
Utilities 8830 530 U 8300 140 F 8440
Supplies 2750 250 U 2500 40 F 2540
Equipment Depreciation 29400 0 None 29400 540 F 29940
Factory rent 8500 400 U 8100 0 None 8100
Property taxes 2600 0 None 2600 0 None 2600
Factory administration 15770 630 F 16400 140 F 16540
Total expenses 144070 1570 U 142500 4380 F 146880